OpenStudy (jazzyfa30):

use the given values to evaluate (if possible)all six trig functions cot theta=-5 sin theta=(squareroot26)/26

OpenStudy (jazzyfa30):

@ganeshie8

ganeshie8 (ganeshie8):

you're given \(\cos\) and \(\sin\) already

ganeshie8 (ganeshie8):

what are the other 4 trig functions u wanto find ?

OpenStudy (jazzyfa30):

sec cot tan csc???

ganeshie8 (ganeshie8):

yup, use below to find them : \(\large \sec = \frac{1}{\cos }\) \(\large \csc = \frac{1}{\sin }\) \(\large \tan = \frac{\sin}{\cos }\) \(\large \cot = \frac{\cos}{\sin }\)

OpenStudy (jazzyfa30):

what now

ganeshie8 (ganeshie8):

use above^

ganeshie8 (ganeshie8):

to get sec, just take reciprocal of cos

OpenStudy (jazzyfa30):

but cotangent not cosine

ganeshie8 (ganeshie8):

Oh,

ganeshie8 (ganeshie8):

we can use another relation : \(\large \cot = \frac{\cos}\sin{}\)

ganeshie8 (ganeshie8):

that gives, \(\large \cos = \cot \times \sin\)

ganeshie8 (ganeshie8):

once u have cos, u can find all others using previous relations..